What force is there between two charges, each with 2.0 x 10-6 C if they are 0.03 m apart?  



1
Expert's answer
2022-05-17T16:04:40-0400
"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{2.0*10^{-6}*2.0*10^{-6}}{0.03^2}=40\:\rm N"